# QADM-Net: Multi-Level Quality-Adaptive Dynamic Network for Reliable Multimodal Classification

### v2

@@ 3.5 Learning
$\mathcal{L}^{task}$ is the classification loss mentioned in Section 3.1. $\Delta=\frac{1}{N}\sum_{i}{\sum_{t}{\Delta_{i}^{t}}}$ where $\Delta_{i}^{t}$ is defined in Equation 7. $\mathcal{L}^{s}$ is the sparsity loss formulated in Equation 9.
4 Experiments
In this section, we conduct experiments on four real-world datasets. The main questions to be verified are highlighted here:
  • Q1: How reliable is QADM-Net in classification?
  • Q2: Are the two proposed dynamic mechanisms effectively improving the model’s reliability?
  • Q3: Is the proposed noise-free prototype confidence estimation (NFCE) more reliable than other confidence estimation methods?
  • Q4: Does QADM-Net learn informative features?
  • Q5: How is QADM-Net’s computation efficiency?
4.1 Experimental Settings
Datasets.
Experiments are conducted on four commonly used multimodal datasets in previous methods Han et al. (2020); Geng et al. (2021); Han et al. (2022); Zhang et al. (2023); Zheng et al. (2023); Zhou et al. (2023); Cao et al. (2024). (1) BRCA is a dataset for classifying breast invasive carcinoma into 5 PAM50 subtypes, containing 875 samples with features from three modalities: mRNA expression, DNA methylation, and miRNA expression. It is available from The Cancer Genome Atlas (TCGA).  [^1] https://www.cancer.gov/aboutnci/organization/ccg/research/structuralgenomics/tcga. (2) ROSMAP contains samples from Alzheimer’s patients (182 samples) and normal control subjects (169 samples) including same modalities as BRCA. A Bennett et al. (2012); De Jager et al. (2018) (3) CUB is an image-text dataset with 11,788 samples comprising 200 categories of birds. Wah et al. (2011) (4) UPMC FOOD101 comprises 90,704 food images from 101 categories and corresponding textual descriptions Wang et al. (2015).
Compared methods.
To demonstrate the improvement in multimodal classification reliability achieved by the proposed QADM-Net, several representative methods in this field are introduced, including multimodal dynamics (MD) Han et al. (2022), multi-level confidence learning (MLCLNet) Zheng et al. (2023), dynamic poly-attention network (DPNET) Zou et al. (2023), enhanced encoding and confidence evaluating framework (CALM) Zhou et al. (2023), quality-aware multimodal fusion (QMF) Zhang et al. (2023), multi-omics data classification network via global and cross-modal feature aggregation (GCFANet) Zheng et al. (2024), and predictive dynamic fusion (PDF) Cao et al. (2024). To demonstrate the advantages of our method over other multimodal dynamic neural network, DynMM Xue and Marculescu (2023) is also included in the experiments. Other dynamic neural networks are excluded from the experiments as they are not applicable to multimodal scenarios.
Evaluation metrics.
@@ 4.1 Experimental Settings
Implementation details.
Noise is added to simulate low-quality data in the experiments, with $\sigma$ representing the noise intensity. We run experiments 20 times to report the mean and standard deviation, following previous work Han et al. (2022). The Adam optimizer and the step learning rate scheduler with an initial learning rate of 1e-4 are employed in the training. More implementation details, including the noise-adding process, are provided in the supplementary material.
4.2 Experimental Results
4.2.1 Classification Performance (Q1)
Following previous works Han et al. (2020); Ma et al. (2021); Zhang et al. (2023); Zhou et al. (2023), we compare the classification performance of different methods under noise-free (Table 1) and Gaussian noise-added (Figure 3) settings. The results demonstrate that: (1) QADM-Net with network dynamics achieves more promising and stable classification performance under different noise intensities compared to its static counterparts. (2) QADM-Net outperforms DynMM. The possible reasons include QADM-Net’s reliable confidence estimation for samples, more robust dynamic depth under the global cross-modality depth normalization, and its capability for dynamic network parameters.
@@ 4.2.1 Classification Performance (Q1)
4.2.2 Ablation study (Q2)
We conduct ablation studies on all four datasets to demonstrate the effectiveness of the Global Confidence Normalized Depth (GCND) and Layer-wise Greedy Parameter (LGP) mechanisms in enhancing the reliability of QADM-Net. To this end, two degraded model versions, “W/O GCND” and “W/O LGP”, are introduced. In “W/O GCND”, the depth of each unimodal model is fixed to the hyperparameter $K$. In “W/O LGP”, the parameters of all network blocks are determined during training instead of dynamically predicted during inference. (i) Effectiveness of GCND. Figure 4 shows the classification accuracy of “W/O GCND” and QADM-Net on four datasets. When $K=1$, the two network structures are identical, resulting in the same accuracy. When $K\geq 2$, QADM-Net’s unimodal networks can dynamically adjust their depth between 1 and $K$. In contrast, “W/O GCND” maintains a fixed depth of $K$, which becomes too deep for a significant portion of easy samples in the dataset, causing overfitting and a notable performance drop. As $K$ increases, the issues with “W/O GCND” become more severe, resulting in a continuous decline in accuracy. (ii) Effectiveness of LGP. Table 2 compares the classification accuracy of “W/O LGP” and QADM-Net on four datasets under different noise intensities. QADM-Net outperforms “W/O LGP” across different noise levels, and its performance degradation as noise intensity increases is significantly less severe than “W/O LGP”. This demonstrates that LGP contributes to both higher classification performance and improved robustness.
We have conducted extensive experiments to show the sample-adaptiveness of QADM-Net, the robustness of GCND to extreme samples, the enhancements in modality and feature quality across layers by GCND and LGP, and so on. Due to space limitations, more experimental results are presented in the supplementary materials.
Table 2: Ablation study on four datasets to demonstrate the effectiveness of LGP, where the best results are in bold.
Dataset Method $\sigma=0$ $\sigma=5$ $\sigma=10$ BRCA W/O LGP 86.2 83.4 80.2 QADM-Net 93.4 92.4 91.5 ROSMAP W/O LGP 85.4 82.6 80.4 QADM-Net 92.2 91.6 90.8 CUB W/O LGP 90.3 86.2 83.6 QADM-Net 97.4 96.7 95.5 FOOD101 W/O LGP 90.2 87.7 85.1 QADM-Net 94.5 93.8 93.0
@@ 4.2.2 Ablation study (Q2)
Table 3: Results of QADM-Net on the FOOD101 dataset using different confidence estimation methods.
@@
QADM-Net (Cls) 93.5 91.7 90.4
QADM-Net (NFCE) 94.5 94.0 93.8
Table 4: Top 5 informative biomarkers that QADM-Net learned from different modalities on BRCA dataset.
Modality Top 5 informative biomarkers DNA ZNF671, KRTAP3-1, ZCCHC8, AGR2, MYT1 mRNA KLK8, MIA, PTX3, SOX11, KRT6B miRNA mir-378, mir-190b, mir-934, mir-21, mir-23b
Table 5: Comparison of ACC and FLOPs per sample of different methods on the FOOD101 dataset, where * are the results under Gaussian noise with $\sigma=5$.
Method GCFANet PDF DynMM QADM-Net ACC (%) 92.9 93.3 89.3 94.5 FLOPs (M) 9.22 3.53 3.71 3.82 ACC (%) 89.4 (-3.5) 90.1 (-3.2) 70.1 (-19.2) 93.8 (-0.7) FLOPs (M) 9.22 (+0) 3.53 (+0) 10.82 (+7.11) 8.94 (+5.12)
4.2.3 Reliability of the Proposed NFCE (Q3)
@@
We conduct experiments on commonly used traditional confidence estimation methods, including MCP, TCP Han et al. (2022), and Energy Liu et al. (2020), all of which assess confidence based on class probabilities obtained by network or classifier. To demonstrate that the proposed noise-free prototype confidence estimation can enhance the reliability of these methods, we replaced their class probability computing approach with the method described in Equation 4, resulting in “MCP-NFCE”, “TCP-NFCE”, and “Energy-NFCE”, respectively. Each method is trained separately on the original and noisy training set with Gaussian noise ($\sigma=5$). Figure 5 reports the confidence estimation of the same clean testing samples provided by two trained versions of each method. The testing results obtained by the same method after training on original and noisy data are expected to be nearly identical since the confidence evaluation is performed on the same test samples. However, the results of the three traditional methods exhibit shifts. Notably, the testing confidence corresponding to the noisy training set is significantly lower than that of the original. The reason is that the classifiers or network structures in these methods overfit the noise distribution, resulting in reduced classification confidence for clean test samples. In contrast, the three “-NFCE” counterparts maintain stable testing confidence for different training noises. Thanks to the noise-free prototypes and the classifier-free design, the confidence of samples can be evaluated more reliably. Furthermore, we substitute the method of Equation 4 in QADM-Net with classifier and yield “QADM-Net (Cls)”. Table 3 compares the classification accuracy of “QADM-Net (Cls)” and original QADM-Net with NFCE under different noise intensities. The results indicate that NFCE provides QADM-Net with greater reliability than classifier-based.
We also conducted experiments showing that noise-free prototypes effectively eliminate data noise and that NFCE achieves multi-level quality estimation. Results are provided in the supplementary material.
4.2.4 Informative features that QADM-Net learns (Q4)

### v3

@@ III-E Learning
$\mathcal{L}^{task}$ is the classification loss mentioned in Sec. III-A. $\Delta=\frac{1}{N}\sum_{i}{\sum_{t}{\Delta_{i}^{t}}}$ where $\Delta_{i}^{t}$ is defined in Eq.(10). $\mathcal{L}^{s}$ is the sparsity loss formulated in Eq.(12).
IV Experiments
In this section, the effectiveness of the proposed Multi-QuAD is evaluated compared with 7 state-of-the-art multimodal reliable classification methods and multimodal dynamic network on four benchmarks. In the following section, we will first introduce the experimental settings in Sec. IV-A. Then, the comparison with SOTA methods, ablation study, discussion on three key component of Multi-QuAD (NFCE, GCND, and LGP), analysis on learned features and computation effectiveness will be presented respectively in Sec. IV-B, IV-C, IV-D, IV-E, IV-F, IV-G, and IV-H.
IV-A Experimental Setups
IV-A1 Datasets
Experiments are conducted on four commonly used multimodal datasets in previous methods [40, 24, 4, 31, 29, 3, 2]. (1) BRCA is a dataset for classifying breast invasive carcinoma into 5 PAM50 subtypes, containing 875 samples with features from three modalities: mRNA expression, DNA methylation, and miRNA expression. It is available from The Cancer Genome Atlas (TCGA)  [^1] https://www.cancer.gov/aboutnci/organization/ccg/research/structuralgenomics/tcga. (2) ROSMAP contains samples from Alzheimer’s patients (182 samples) and normal control subjects (169 samples), including the same modalities as BRCA. [59, 60] (3) CUB is an image-text dataset with 11,788 samples comprising 200 categories of birds. [61] (4) UPMC FOOD101 comprises 90,704 food images from 101 categories and corresponding textual descriptions [62].
IV-A2 Compared methods
To demonstrate the improvement in multimodal classification reliability achieved by Multi-QuAD, several representative methods in this field are introduced, including multimodal dynamics (MD) [4], multi-level confidence learning (MLCLNet) [29], dynamic poly-attention network (DPNET) [28], enhanced encoding and confidence evaluating framework (CALM) [3], quality-aware multimodal fusion (QMF) [31], multi-omics data classification network via global and cross-modal feature aggregation (GCFANet) [1], and predictive dynamic fusion (PDF) [2]. To demonstrate the advantages of our method over other multimodal dynamic neural networks, DynMM [56] is also included in the experiments. Other dynamic neural networks are excluded from the experiments as they are not applicable to multimodal scenarios.
IV-A3 Evaluation metrics
@@
IV-A4 Implementation details
Noise is added to simulate low-quality data in the experiments, with $\sigma$ representing the noise intensity. We run experiments 20 times to report the mean and standard deviation, following previous work [4]. The Adam optimizer and the step learning rate scheduler with an initial learning rate of 1e-4 are employed in the training. More details are provided in the Supplemental Materials.
IV-B Comparison With State-of-The-Art Methods
Following previous works [40, 63, 31, 3], we compare the classification performance of different methods under noise-free and Gaussian noise-added settings, as shown in Table I and Fig. 4, respectively.
(1) Comparison with state-of-the-art reliable multimodal classification methods. Results in Table I show that the proposed Multi-QuAD outperforms SOTA reliable multimodal classification methods such as CALM, GCFANet, and PDF by a substantial margin under the noise-free setting on four datasets. On BRCA, ROSMAP, CUB, and FOOD101, Multi-QuAD achieves ACC gains of 4.8%, 5.9%, 4.4%, and 1.2% over the leading approach, respectively. The possible reason is that Multi-QuAD’s dynamic network capability enables it to learn distinct mapping relationships for different samples, thereby enhancing the precision of the predictions. Under noise setting (Fig. 4), Multi-QuAD exhibits more stable performance compared to SOTA methods, with a larger ACC improvement than under noise-free conditions. This indicates that the reliable quality estimation method, the sample-specific network depth, and sample-specific parameters endow Multi-QuAD with a more robust performance.
(2) Comparison with multimodal dynamic neural network. Multi-QuAD outperforms DynMM across four datasets under the noise-free setting as shown in Table I. This is primarily because Multi-QuAD’s dynamic parameter mechanism provides stronger representation capability. While under Gaussian noise (Fig. 4), Multi-QuAD’s more reliable quality estimation method and dynamic depth mechanism make its performance much more stable than DynMM. DynMM suffers significant performance degradation as noise intensity $\sigma$ increases. This is likely because it adjusts excessively deep network layers for hard samples, leading to overfitting and unreliable results.
TABLE II: Ablation study on four datasets to demonstrate the effectiveness of LGP, where the best results are in bold.
Dataset Method $\sigma=0$ $\sigma=5$ $\sigma=10$
BRCA W/O LGP 86.2 83.4 80.2
Multi-QuAD 93.4 92.4 91.5
ROSMAP W/O LGP 85.4 82.6 80.4
Multi-QuAD 92.2 91.6 90.8
CUB W/O LGP 90.3 86.2 83.6
Multi-QuAD 97.4 96.7 95.5
FOOD101 W/O LGP 90.2 87.7 85.1
Multi-QuAD 94.5 93.8 93.0
IV-C Ablation study
Ablation studies are conducted on all four datasets to demonstrate the effectiveness of the Global Confidence Normalized Depth (GCND) and Layer-wise Greedy Parameter (LGP) mechanisms in enhancing the reliability of Multi-QuAD. To this end, two degraded model versions, “W/O GCND” and “W/O LGP”, are introduced. In “W/O GCND”, the depth of each unimodal model is fixed to the hyperparameter $K$. In “W/O LGP”, the parameters of all network blocks are determined during training instead of being dynamically predicted during inference. (i) Effectiveness of GCND. Fig. 5 shows the classification accuracy of “W/O GCND” and Multi-QuAD on four datasets. When $K=1$, the two network structures are identical, resulting in the same accuracy. When $K\geq 2$, Multi-QuAD’s unimodal networks can dynamically adjust their depth between 1 and $K$. In contrast, “W/O GCND” maintains a fixed depth of $K$, which becomes too deep for a significant portion of easy samples in the dataset, causing overfitting and a notable performance drop. As $K$ increases, the issues with “W/O GCND” become more severe, resulting in a continuous decline in accuracy. (ii) Effectiveness of LGP. Table II compares the classification accuracy of “W/O LGP” and Multi-QuAD on four datasets under different noise intensities. Multi-QuAD outperforms “W/O LGP” across different noise levels, and its performance degradation as noise intensity increases is significantly less severe than “W/O LGP”. This demonstrates that LGP contributes to both higher classification performance and improved robustness.
@@ IV-C Ablation study
(a) Results on the BRCA dataset.
(b) Results on the UPMC FOOD101 dataset.
Figure 7: The probability of noise-free prototypes in the noise distribution $P_{\mathcal{N}}(\omega)$ at different optimization epochs.
TABLE III: Results of Multi-QuAD on the FOOD101 dataset using different confidence estimation methods.
@@
Multi-QuAD (Cls) 93.5 91.7 90.4
Multi-QuAD (NFCE) 94.5 94.0 93.8
IV-D Discussion on NFCE
IV-D1 The Reliability of NFCE
@@
We conduct experiments on commonly used traditional confidence estimation methods, including MCP, TCP [4], and Energy [45], all of which assess confidence based on class probabilities obtained by network or classifier. To demonstrate that the proposed noise-free prototype confidence estimation can enhance the reliability of these methods, we replaced their class probability computing approach with the method described in Eq.(4), resulting in “MCP-NFCE”, “TCP-NFCE”, and “Energy-NFCE”, respectively. Each method is trained separately on the original and noisy training set with Gaussian noise ($\sigma=5$). Fig. 6 reports the confidence estimation of the same clean testing samples provided by two trained versions of each method. The testing results obtained by the same method after training on original and noisy data are expected to be nearly identical since the confidence evaluation is performed on the same test samples. However, the results of the three traditional methods exhibit shifts. Notably, the testing confidence corresponding to the noisy training set is significantly lower than that of the original. The reason is that the classifiers or network structures in these methods overfit the noise distribution, resulting in reduced classification confidence for clean test samples. In contrast, the three “-NFCE” counterparts maintain stable testing confidence for different training noises. Thanks to the noise-free prototypes and the classifier-free design, the confidence of samples can be evaluated more reliably.
Furthermore, we substitute the method of Eq.(4) in Multi-QuAD with classifier and yield “Multi-QuAD (Cls)”. Table III compares the classification accuracy of “Multi-QuAD (Cls)” and original Multi-QuAD with NFCE under different noise intensities. The results indicate that NFCE provides Multi-QuAD with greater reliability than classifier-based.
TABLE IV: Comparison of classification accuracy under different noise intensities for Multi-QuAD and its “depth-plus-one version” and “depth-minus-one version.”
Dataset Method $\sigma=0$ $\sigma=5$ $\sigma=10$
BRCA Multi-QuAD(-1) 92.8 91.4 90.2
Multi-QuAD(+1) 92.9 91.6 90.1
Multi-QuAD 93.7 92.4 91.5
ROSMAP Multi-QuAD(-1) 91.6 90.8 90.1
Multi-QuAD(+1) 91.5 90.8 90.2
Multi-QuAD 92.1 91.6 90.8
CUB Multi-QuAD(-1) 96.4 95.7 95.0
Multi-QuAD(+1) 96.5 95.7 94.9
Multi-QuAD 97.3 96.7 95.5
FOOD101 Multi-QuAD(-1) 93.5 92.9 92.3
Multi-QuAD(+1) 93.5 92.8 92.2
Multi-QuAD 94.4 93.7 93.0
IV-D2 Noise Removal of the Noise-Free Prototypes (The Effectiveness of $\mathcal{L}^{rob}$)
To verify that the noise is effectively removed in prototypes during optimization via $\mathcal{L}^{rob}$ formulated in Eq.(3), noise with $\sigma=5$ is added to all modalities of the datasets. At each optimization epoch, we report the probability of the prototype corresponding to each class and modality within the noise distribution: $P_{\mathcal{N}}(\omega)=\{\{P_{\mathcal{N}}(\omega^{c,m})\}_{c=1}^{C}\}_{m=1}^{M}$. $P_{\mathcal{N}}$ is the probability density function of the added noise. The results in Fig. 7a and 7b shows that as the optimization progresses, the probabilities $P_{\mathcal{N}}(\omega)$ quickly converge to nearly 0 after a few epochs, indicating that the noise in the prototypes is effectively removed using $\mathcal{L}^{rob}$.
IV-D3 Multi-level quality estimation
Extensive experiments are conducted to verify NFCE’s capability to provide modality-level and feature-level quality estimation. (i) Modality level. As shown in Fig. 8a, noise is added to the text modality of half the UPMC FOOD101 dataset samples, and the quality distributions of both modalities calculated using $\delta_{v}$ (Eq.(6)) in these two halves are reported. When the noise intensity $\sigma=0$, the quality distributions of the two halves of the text modality samples are nearly identical. As noise intensity in text modality increases, the quality distribution of the noisy text data becomes increasingly skewed to the left, while the unaffected image modality maintains consistent quality between the two halves of the data. This indicates that NFCE can efficiently sense changes in modality quality. (ii) Feature level. Fig. 8b visualizes the quality of the noisy half of text modality, with each element in each row of the heatmap representing the quality of each feature in a sample’s feature vector. The distinct values in each column of the heatmap indicate that NFCE can perceive the varying qualities of different feature values among samples calculated using $\delta_{f}$ (Eq.(7)). As the noise intensity increases, the feature-level quality of the samples decreases, and the overall color of the heatmap becomes lighter. This indicates that NFCE can perceive the changes in feature-level quality brought about by noise intensity.
(a) The quality distributions of clean and noisy halves of the data in different modalities on the UPMC FOOD101 dataset.
(b) The feature-level quality of the noisy data on the UPMC FOOD101 dataset under different noise intensities.
Figure 8: Results of the modality-level and feature-level quality estimated by NFCE.
IV-E Discussion on the GCND
IV-E1 GCND Adjusts Optimal Network Depth For Samples
Two modified versions of Multi-QuAD terms “Multi-QuAD(-1)” and “Multi-QuAD(+1)” are introduced to verify whether GCND adjusts the optimal depth for samples. “Multi-QuAD(-1)” employs network depths that are one layer shallower than Multi-QuAD. Specifically, the network depths of “Multi-QuAD(-1)” for sample $x_{i}$ is $\textbf{D}_{i}^{\prime}=\{max(\textbf{D}_{i}^{m}-1,1)\}_{m=1}^{M}$, where $\textbf{D}_{i}=\{\textbf{D}_{i}^{m}\}_{m=1}^{M}$ is the depth adjusted by Multi-QuAD. “Multi-QuAD(+1)” employs network depths that are one layer deeper than Multi-QuAD, i.e., $\textbf{D}_{i}^{\prime}=\{\textbf{D}_{i}^{m}+1\}_{m=1}^{M}$. We add noise with different intensities $\sigma$ and compare the classification accuracy of “Multi-QuAD(-1)”, “Multi-QuAD(+1)”, and Multi-QuAD on all datasets. Results in Table IV demonstrate that both “Multi-QuAD(-1)” and “Multi-QuAD(+1)” consistently underperform compared to Multi-QuAD across all four datasets and various noise intensities. This indicates that Multi-QuAD adjusts the most reliable depths for samples.
Figure 9: Visualization of the parameter values of different samples on UMPC FOOD101 dataset under different noise intensities. For clarity, the first 10 elements of the parameters are shown.
TABLE V: Classification accuracy of ICD and GCND on four datasets under different noise intensities.
Dataset Method $\sigma=0$ $\sigma=5$ $\sigma=10$
BRCA ICD 92.0 89.4 87.2
GCND 93.4 92.4 91.5
ROSMAP ICD 90.4 88.2 86.5
GCND 92.2 91.6 90.8
CUB ICD 94.6 92.1 89.4
GCND 97.4 96.7 95.5
FOOD101 ICD 92.1 90.5 88.6
GCND 94.5 93.8 93.0
IV-E2 GCND Exhibits Stronger Reliability Than Conventional Dynamic Depth Method
Existing dynamic depth methods typically rely on intermediate modules to determine network depth, such as the gating network in DynMM [56] and the intermediate classifier in MSDNet [50]. The design of these intermediate modules prevents the network depth from being determined before sample input, making the methods prone to excessive depth for difficult samples during inference and thus affecting model reliability. In contrast, GCND enhances the reliability of dynamic depth by mitigating the adverse effects of extreme modality samples, as introduced in Sec. III-C. To validate this, we implement the intermediate-classifier-based dynamic depth termed “ICD” following that in MSDNet, and replace GCND in Multi-QuAD with it. The classification accuracy of the two implementations of dynamic depth under different noise intensities is illustrated in Table V. It is evident that GCND enables Multi-QuAD to achieve higher and more stable classification accuracy than ICD under varying noise levels, demonstrating its stronger reliability.
IV-E3 Analysis of Hyperparameter $K$
There is one hyperparameter $K$ in GCND (mentioned in Sec. III-C). The sensitivity analysis is conducted by varying $K$ in the range $[1,10]$. The result in the purple curve of Fig. 5 shows that Multi-QuAD is relatively insensitive to $K$ in all datasets as $K$ is greater than or equal to 3. Due to space limitation, more discussions on GCND can be found in the Supplemental Materials.
Figure 10: Visualization of the element values of the hidden state and cell state of the LSTM Unit at the first layer. The result on UPMC FOOD101 dataset is shown, the noisy half of the samples are marked with red box.
(a) The results on the BRCA dataset.
(b) The results on the ROSMAP dataset.
Figure 11: The t-SNE visualization results of the output feature vectors of different modalities enhanced by Multi-QuAD on the BRCA and ROSMAP datasets. Clusters of different colors represent the features of samples from different classes.
IV-F Discussion on the LGP
IV-F1 Parameters Adaptability to Feature-Level Quality
To demonstrate the adaptability of parameters provided by LGP under feature-level quality variation, we visualize the parameters from the same quality-enhancing network block of Multi-QuAD for different samples under noise. Fig. 9 shows the parameter values of 10 randomly chosen samples of the noisy text modality on the UPMC FOOD101 dataset. By comparing the differences between the rows of the heatmaps, one can observe that LGP predicts fairly consistent yet distinct network parameters for clean samples ($\sigma=0$). However, as the intensity of noise increases, introducing more chaotic features into the feature vectors, the network parameters predicted by LGP for different samples become increasingly varied. This demonstrates that LGP can perceive the quality variations at the feature level within the feature vectors of each sample and dynamically adjust the parameters.
IV-F2 Effectiveness of the LSTM Unit (LU)
To further illustrate the mechanism of LGP, we visualize the hidden state and cell state of the LSTM Unit (LU) under noise. Fig. 10 shows the results at the first layer on the UPMC FOOD101 dataset. Under noise-free conditions, the hidden state and cell state across samples are rather consistent. After the noise ($\sigma=10$) is added to half of the samples, the feature-level quality of the affected half of the samples becomes irregular and varied, and the corresponding half of the hidden state and cell state (marked with red boxes) also exhibit cross-sample variation. This demonstrates that LU can perceive changes in feature values across samples, thereby ensuring reliable parameter prediction. Due to space limitation, more discussions on LGP can be found in the Supplemental Materials.
IV-G Informative features that Multi-QuAD learns